# Provably Faster Gradient Descent via Long Steps

### v1

@@ 2.1 Convergence Guarantees for Straightforward Stepsize Patterns
Note straightforwardness provides no guarantees on the intermediate objective values of at iterations $k=1,2,\dots t-1$. We show in our theorem below descent at every $t$th iteration. To allow for some numerical flexibility, we say a pattern is $\epsilon$-straightforward if for some $\Delta\in(0,1/2]$, all $\delta\in[0,LD^{2}\Delta]$ have value function bounded by $p_{L,D}(\delta)\leq\delta-\frac{\sum_{i=0}^{t-1}(h_{i}-\epsilon)}{LD^{2}}\delta^{2}$.
@@
Consider any $L$-smooth, convex $f$. If $h=(h_{0},\dots,h_{t-1})$ is $\epsilon$-straightforward with parameter $\Delta\in(0,1/2]$, then gradient descent (1.3) has, for any $s,t\in\mathbb{N}$, after $T=st$ gradient steps
@@
where $\bar{s}=\left\lceil\frac{\log\left(\frac{f(x_{0})-f(x_{\star})}{LD^{2}\Delta}\right)}{\log(1-\sum_{i=0}^{t-1}(h_{i}-\epsilon)\Delta)}\right\rceil$. In particular, suppressing lower-order terms, this rate is
@@
We begin by showing $\epsilon\geq 0$-straightforwardness implies for any $s\in\mathbb{N}$, the recurrence relation
@@
Consider first $s=0$. The first case of (2.4) follows from the definition of straightforwardness as $\delta_{t}\leq p_{L,D}(\delta_{0})\leq\delta_{0}-\frac{\sum_{i=0}^{t-1}(h_{i}-\epsilon)}{LD^{2}}\delta_{0}^{2}$. The second case of (2.4) follows from observing that $p_{L,D}(\cdot)$ is concave: Since $p_{L,D}(0)=0$ and $p_{L,D}\left(LD^{2}\Delta\right)\leq LD^{2}\left(\Delta-\sum_{i=0}^{t-1}(h_{i}-\epsilon)\Delta^{2}\right)$, every $\delta_{0}>LD^{2}\Delta$ must have
@@
Next, we show this recurrence implies the claimed convergence rate (2.3). Suppose first, $\delta_{st}$ is larger than $LD^{2}\Delta$. Then applying the pattern $h$ contracts the objective gap [^4] Interestingly, this contraction factor is independent of $L,D$. As a result, problem conditioning plays a minimal role in this initial phase of convergence. Instead, only $h$ and the associated straightforwardness parameter $\Delta$ matter., inductively giving $\delta_{st}\leq(1-\sum_{i=0}^{t-1}(h_{i}-\epsilon)\Delta)^{s}\delta_{0}$. After at most $\bar{s}$ executions of the stepsize pattern, one must have $\delta_{st}\leq LD^{2}\Delta$. Afterward, for any $s>\bar{s}$, the objective gap decreases by at least
@@
Solving this recurrence with the initial condition $\delta_{\bar{s}t}\leq LD^{2}\Delta$ gives the claimed sublinear rate
$\delta_{st}\leq\frac{LD^{2}}{\sum_{i=0}^{t-1}(h_{i}-\epsilon)(s-\bar{s})+\frac{1}{\Delta}}\ .\qed$
@@ 2.2 Faster Convergence for Straightforward Patterns given Growth Bounds
Strong convexity implies this condition with $q=2$ and leads gradient descent with constant $h=(1,1,\dots)$ to converge at a rate of $O((1-\mu/L)^{T})$. When $q>2$, improved sublinear guarantees of $O((L/\mu^{2/q}T)^{q/(q-2)})$ follow. Below we show that any straightforward stepsize pattern enjoys the same convergence improvements gaining an additional factor of $\mathrm{avg}(h)$.
@@
Consider any $L$-smooth, convex objective $f$ satisfying (2.5). If $h=(h_{0},\dots,h_{t-1})$ is $\epsilon$-straightforward with parameter $\Delta\in(0,1/2]$, then gradient descent (1.3) has, for any $s,t\in\mathbb{N}$ with $s>\bar{s}$, after $T=st$ gradient steps
$f(x_{T})-f(x_{\star})\leq\begin{cases}q\left(\dfrac{L}{\mu^{2/q}(q-2)(\mathrm{avg}(h)-\epsilon)(T-\bar{s}t)}\right)^{q/(q-2)}&\text{ if }q>2\\ \left(1-\frac{\mu(\mathrm{avg}(h)-\epsilon)t}{2L}\right)^{(s-\bar{s})}LD^{2}\Delta&\text{ if }q=2\ .\\ \end{cases}$
@@
$\delta_{(s+1)t}\leq\begin{cases}\delta_{st}-\frac{\mu^{2/q}\sum_{i=0}^{t-1}(h_{i}-\epsilon)}{Lp^{2/q}}\delta_{st}^{2-2/q}&\text{ if }\delta_{st}\leq LD_{0}^{2}\Delta\\ \left(1-\sum_{i=0}^{t-1}(h_{i}-\epsilon)\Delta\right)\delta_{0}&\text{ if }\delta_{st}>LD_{0}^{2}\Delta\ .\end{cases}$
The analysis of initial linear convergence is unchanged. Once $\delta_{0}\leq LD_{0}^{2}\Delta$ (after at most $\bar{s}$ applications of the straightforward stepsize pattern), convergence occurs based on the recurrence relation $\delta_{(s+1)t}\leq\delta_{st}-\frac{\mu^{2/q}\sum_{i=0}^{t-1}h_{i}}{Lq^{2/q}}\delta_{st}^{2-2/q}$. Solving this modified recurrence relation (for example, see [23, Lemma A.1] for this calculation) shows for any $s>\bar{s}$,
$\delta_{st}\leq\begin{cases}q\left(\dfrac{L}{\mu^{2/q}(q-2)\sum_{i=0}^{t-1}(h_{i}-\epsilon)(s-\bar{s})}\right)^{q/(q-2)}&\text{ if }q>2\\ \left(1-\frac{\mu\sum_{i=0}^{t-1}(h_{i}-\epsilon)}{2L}\right)^{(s-\bar{s})}LD_{0}^{2}\Delta&\text{ if }q=2\ .\\ \end{cases}\qed$
@@ 3 Certificates of Straightforwardness
All that remains is to show how one can certify the straightforwardness of a stepsize pattern. We do this in two steps. First, Section 3.1 shows that $p_{L,D}(\delta)$ is upper bounded by an SDP minimization problem using PEP techniques. Hence straightforwardness is implied by showing the SDP corresponding to each $\delta\in[0,LD^{2}\Delta]$ has a sufficiently good feasible solution. Second, Section 3.2 shows that another semidefinite programming feasibility problem can certify that such an interval of solutions exists.
@@ 3.1 The Worst-Case Value Function is Upper Bounded by an SDP
We first reformulate the infinite-dimensional problem (2.1) as a finite-dimensional nonconvex quadratically constrained quadratic problem (see (3.1)), relax that formulation into a SDP (see (3.2)), and then upper bound the SDP by its dual problem (see (3.3)). This process has been detailed by [1, 4, 12] and is carried out in our setting below (following the notations of [12]).
Step 1: A QCQP reformulation. First, using the interpolation theorem of [2], $p_{L,D}(\delta)$ equals the following finite-dimensional nonconvex problem optimizing over all possible objective values $f_{k}$ and gradients $g_{k}$ at the points $x_{k}$ with $k\in I_{t}^{\star}:=\{\star,0,1,\dots t\}$
@@
Step 2: An SDP relaxation. Second, we relax the nonconvex problem (3.1) to the following SDP. Following the notation of [12], define
@@
where $x\odot y=\frac{1}{2}(xy^{T}+yx^{T})$ denotes the symmetric outer product. This notation is defined so that $g_{j}^{T}(x_{i}-x_{j})=\mathrm{Tr}GA_{i,j}(h)$, $\|x_{i}-x_{j}\|^{2}_{2}=\mathrm{Tr}GB_{i,j}(h)$, and $\|g_{i}-g_{j}\|^{2}_{2}=\mathrm{Tr}GC_{i,j}$ for any $i,j\in I_{t}^{\star}$. Then the QCQP formulation (3.1) can be relaxed [^5] Under an additional rank condition (that the problem dimension $n$ exceeds $t+2$), the QCQP problem (3.1) and SDP (3.2) are actually equivalent. However, this is not needed for our analysis, so we make no such assumption. to
$p_{L,D}(\delta)\leq$$\begin{cases}\max_{F,G}&F\mathbf{f}_{t}\\ \textrm{s.t.}&Fa_{i,j}+\mathrm{Tr}GA_{i,j}(h)+\frac{1}{2L}\mathrm{Tr}GC_{i,j}\leq 0,\quad i,j\in I_{t}^{\star}:i\neq j\\ &-G\preceq 0\\ &\mathrm{Tr}GB_{0,\star}\leq D^{2}\\ &F\mathbf{f}_{0}\leq\delta\end{cases}$ (3.2)
with decision variables $F\in\mathbb{R}^{1\times(t+1)}$ and $G\in\mathbb{R}^{(t+2)\times(t+2)}$.
Step 3: The upper bounding dual SDP. Third, we note the maximization SDP (3.2) is bounded above by its dual minimization SDP by weak duality (note strong duality is not assumed):
$p_{L,D}(\delta)\leq\begin{cases}\min_{\lambda,v,w,Z}&D^{2}v+\delta w\\ \textrm{s.t.}&\sum_{i,j\in I_{t}^{\star}:i\neq j}\lambda_{i,j}a_{i,j}=a_{\star,t}-wa_{\star,0}\\ &vB_{0,\star}+\sum_{i,j\in I_{t}^{\star}:i\neq j}\lambda_{i,j}\left(A_{i,j}(h)+\frac{1}{2L}C_{i,j}\right)=Z\\ &Z\succeq 0\\ &v,w\geq 0,\;\lambda_{i,j}\geq 0,\quad i,j\in I_{t}^{\star}:i\neq j\ .\end{cases}$ (3.3)
@@ 3.2 An SDP Feasibility Certificate that implies Straightforwardness
where $m_{h}\colon\mathbb{R}^{(t+2)\times(t+2)}\rightarrow\mathbb{R}^{t+1}$ and $M_{h}\colon\mathbb{R}^{(t+2)\times(t+2)}\rightarrow\mathbb{R}^{(t+1)\times(t+1)}$ are linear functions. Certifying $p(\delta)\leq\delta-\sum_{i=0}^{t-1}(h_{i}-\epsilon)\delta^{2}$ for fixed $\delta$ then follows by showing the following spectral set is nonempty
@@
$\begin{bmatrix}\sum_{i=0}^{t-1}(h_{i}+\epsilon)\delta^{2}&\delta m_{h}(\gamma)^{T}\\ \delta m_{h}(\gamma)&M_{h}(\lambda)+\delta M(\gamma)\end{bmatrix}\succeq 0\ .$
@@
$\begin{bmatrix}\sum_{i=0}^{t-1}(h_{i}+\epsilon)&m_{h}(\gamma)^{T}\\ m_{h}(\gamma)&M(\lambda)+\delta M_{h}(\gamma)\end{bmatrix}\succeq 0\ .$
@@
3.3 Certificates of Straightforwardness Proving Faster Rates in Table 1

### v5

@@ 2.1 Convergence Guarantees for Straightforward Stepsize Patterns
Note straightforwardness provides no guarantees on the intermediate objective values at iterations $k=1,2,\dots t-1$. We show in our theorem below descent at every $t$th iteration. To allow for some numerical flexibility, we say a pattern is $\epsilon$-straightforward if for some $\Delta\in(0,1/2]$, all $\delta\in[0,LD^{2}\Delta]$ have value function bounded by $p_{L,D}(\delta)\leq\delta-\frac{\sum_{i=0}^{t-1}(h_{i}-\epsilon)}{LD^{2}}\delta^{2}$.
@@
Consider any $L$-smooth, convex $f$. If $h=(h_{0},\dots,h_{t-1})$ is $\epsilon$-straightforward with parameter $\Delta\in(0,1/2]$, then gradient descent (1.3) has, for any $s\in\mathbb{N}$, after $T=st$ gradient steps
@@
where $\bar{s}=\left\lceil\frac{\log\left(\frac{f(x_{0})-f(x_{\star})}{LD^{2}\Delta}\right)}{\log(1-\sum_{i=0}^{t-1}(h_{i}-\epsilon)\Delta)}\right\rceil$. In particular, suppressing lower-order terms, this guarantee is
@@
For any $\epsilon\geq 0$, we begin by showing $\epsilon$-straightforwardness implies that for any $s\in\mathbb{N}$, the recurrence relation
@@
Consider first $s=0$. The first case of (2.4) follows from the definition of straightforwardness as $\delta_{t}\leq p_{L,D}(\delta_{0})\leq\delta_{0}-\frac{\sum_{i=0}^{t-1}(h_{i}-\epsilon)}{LD^{2}}\delta_{0}^{2}$. For the second case of (2.4), first note that $p_{L,D}(\cdot)$ is concave by Lemma 2.1. Then since $p_{L,D}(0)=0$ and $p_{L,D}\left(LD^{2}\Delta\right)\leq LD^{2}\left(\Delta-\sum_{i=0}^{t-1}(h_{i}-\epsilon)\Delta^{2}\right)$, every $\delta_{0}>LD^{2}\Delta$ must have
@@
Next, we show this recurrence implies the claimed convergence guarantee (2.3). Suppose first, $\delta_{st}$ is larger than $LD^{2}\Delta$. Then applying the pattern $h$ contracts the objective gap [^4] Interestingly, this contraction factor is independent of $L,D$. As a result, problem conditioning plays a minimal role in this initial phase of convergence. Instead, only $h$ and the associated straightforwardness parameter $\Delta$ matter., inductively giving $\delta_{st}\leq(1-\sum_{i=0}^{t-1}(h_{i}-\epsilon)\Delta)^{s}\delta_{0}$. After at most $\bar{s}$ executions of the stepsize pattern, one must have $\delta_{st}\leq LD^{2}\Delta$. Afterward, for any $s>\bar{s}$, the objective gap decreases by at least
@@
Solving this recurrence with the initial condition $\delta_{\bar{s}t}\leq LD^{2}\Delta$ gives
$\delta_{st}\leq\frac{LD^{2}}{\sum_{i=0}^{t-1}(h_{i}-\epsilon)(s-\bar{s})+\frac{1}{\Delta}}\ .$
@@ 2.2 Faster Convergence for Straightforward Patterns given Growth Bounds
Strong convexity implies this condition with $q=2$ and leads gradient descent with constant $h=(1,1,\dots)$ to converge at a rate of $O((1-\mu/L)^{T})$. When $q>2$, improved sublinear guarantees of $O((L/\mu^{2/q}T)^{q/(q-2)})$ follow. Below, we show that any straightforward stepsize pattern enjoys the same convergence improvements, gaining a similar factor of $\mathrm{avg}(h)$.
@@
Consider any $L$-smooth, convex objective $f$ satisfying (2.5). If $h=(h_{0},\dots,h_{t-1})$ is $\epsilon$-straightforward with parameter $\Delta\in(0,1/2]$, then gradient descent (1.3) has, for any $T=st$
$f(x_{T})-f(x_{\star})\leq\left(1-\min\left\{\Delta,\ \frac{\mu}{2L}\right\}\sum_{i=0}^{t-1}(h_{i}-\epsilon)\right)^{s}(f(x_{0})-f(x_{\star}))$
when $q=2$ and when $q>2$ and $s>\bar{s}:=\left\lceil\frac{\log\left(\frac{f(x_{0})-f(x_{\star})\mu^{2/(q-2)}}{q^{2/(q-2)}(L\Delta)^{q/(q-2)}}\right)}{\log(1-\sum_{i=0}^{t-1}(h_{i}-\epsilon)\Delta)}\right\rceil$ has
$f(x_{T})-f(x_{\star})\leq q\left(\dfrac{L}{\mu^{2/q}(q-2)(\mathrm{avg}(h)-\epsilon)((s-\bar{s})t)}\right)^{q/(q-2)}\ .$
@@
$\delta_{(s+1)t}\leq\begin{cases}\delta_{st}-\frac{\mu^{2/q}\sum_{i=0}^{t-1}(h_{i}-\epsilon)}{Lq^{2/q}}\delta_{st}^{2-2/q}&\text{ if }\delta_{st}\leq LD^{2}_{st}\Delta\\ \left(1-\sum_{i=0}^{t-1}(h_{i}-\epsilon)\Delta\right)\delta_{st}&\text{ if }\delta_{st}>LD_{st}^{2}\Delta\ .\end{cases}$
One can bound this recurrence relation by the maximum of the two cases. If $q=2$, both cases give a contraction and so
$\delta_{st}\leq\left(1-\min\left\{\Delta,\ \frac{\mu}{2L}\right\}\sum_{i=0}^{t-1}(h_{i}-\epsilon)\right)^{s}\delta_{0}\ .$
If $q>2$, one has
$\delta_{(s+1)t}\leq\max\left\{\delta_{st}-\frac{\mu^{2/q}\sum_{i=0}^{t-1}(h_{i}-\epsilon)}{Lq^{2/q}}\delta_{st}^{2-2/q},\ \left(1-\sum_{i=0}^{t-1}(h_{i}-\epsilon)\Delta\right)\delta_{st}\right\}\ .$
Solving this recurrence gives linear convergence until the objective gap is less than $\left(\frac{q^{2/q}L\Delta}{\mu^{2/q}}\right)^{\frac{q}{q-2}}$ is reached, which must occur by iteration $\bar{s}$, and then afterwards gives a sublinear convergence rate (for example, see [34, Lemma A.1] for this calculation), giving the claim. ∎
Note that the above bound for the case of $q>2$ is independent of $D$. Asymptotically, this is reasonable as better bounds on distance to optimal are eventually available, namely $(\frac{q}{\mu}\delta_{k})^{1/q}$. One could give a bound depending on $D=D_{0}$ early onby using the stronger bound $D_{k}\leq\min\{(\frac{q}{\mu}\delta_{k})^{1/q},D_{0}\}$. However, once the gap is small enough for the first term to dominate, the bound will again be independent of $D_{0}$.
@@ 3 Certificates of Straightforwardness
All that remains is to show how one can certify the straightforwardness of a stepsize pattern. We do this in two steps. First, Section 3.1 shows that $p_{L,D}(\delta)$ is upper bounded by an SDP minimization problem using previously developed PEP techniques. Hence straightforwardness is implied by showing the SDP corresponding to each $\delta\in[0,LD^{2}\Delta]$ has a sufficiently good feasible solution. Second, Section 3.2 shows that another semidefinite programming feasibility problem can certify that such an interval of solutions exists.
@@ 3.1 The Worst-Case Value Function is Upper Bounded by an SDP
We first reformulate the infinite-dimensional problem (2.1) as a finite-dimensional nonconvex quadratically constrained quadratic problem (see (3.1)), relax that formulation into an SDP (see (3.2)), and then upper bound the SDP by its dual problem (see (3.3)). This process was been developed by [1, 2, 3] and is carried out in our particular setting below.
Step 1: A QCQP reformulation. First, as proposed by Drori and Teboulle [1], one can discretize the infinite-dimensional problem defining $p_{L,D}(\delta)$ over all possible objective values $f_{k}$ and gradients $g_{k}$ at the points $x_{k}$ with $k\in I_{t}^{\star}:=\{\star,0,1,\dots t\}$ as done below. Using the interpolation theorem of Taylor et al. [2], this is an exact reformulation rather than a relaxation, giving
@@
Step 2: An SDP relaxation. Second, one can relax the nonconvex problem (3.1) to the following SDP as done in [2, 3, 15]. We follow the particular notational choices of [13], which recently considered globally, numerically optimizing gradient descent’s stepsizes with fixed total iterations, defining
@@
where $x\odot y=\frac{1}{2}(xy^{T}+yx^{T})$ denotes the symmetric outer product. This notation is defined so that $g_{j}^{T}(x_{i}-x_{j})=\mathrm{Tr}GA_{i,j}(h)$, $\|x_{i}-x_{j}\|^{2}_{2}=\mathrm{Tr}GB_{i,j}(h)$, and $\|g_{i}-g_{j}\|^{2}_{2}=\mathrm{Tr}GC_{i,j}$ for any $i,j\in I_{t}^{\star}$. Then the QCQP formulation (3.1) can be relaxed to
$p_{L,D}(\delta)\leq$$\begin{cases}\max_{F,G}&F\mathbf{f}_{t}\\ \textrm{s.t.}&Fa_{i,j}+\mathrm{Tr}GA_{i,j}(h)+\frac{1}{2L}\mathrm{Tr}GC_{i,j}\leq 0,\quad\forall i,j\in I_{t}^{\star}:i\neq j\\ &-G\preceq 0\\ &\mathrm{Tr}GB_{0,\star}\leq D^{2}\\ &F\mathbf{f}_{0}\leq\delta\ .\end{cases}$ (3.2)
Equality holds above if one adds the constraint that $\mathrm{rank\ }G\leq n$ where $n$ is the dimension of problem instances considered when defining $p_{L,D}$. This constraint is vacuously true if the considered problem dimension $n$ exceeds $t+2$. Consequently, the QCQP (3.1) and the SDP (3.2) are equivalent, provided one allows sufficiently high-dimensional objectives, but may differ if one restricts to finding the worst-case over lower-dimensional problem instances. However, this equality is not needed for our analysis, so we make no such assumption.
Step 3: The upper bounding dual SDP. Third, note the maximization SDP (3.2) is bounded above by its dual minimization SDP by weak duality, giving
$p_{L,D}(\delta)\leq\begin{cases}\min_{\lambda,v,w,Z}&D^{2}v+\delta w\\ \textrm{s.t.}&\sum_{i,j\in I_{t}^{\star}:i\neq j}\lambda_{i,j}a_{i,j}=a_{\star,t}-wa_{\star,0}\\ &vB_{0,\star}+\sum_{i,j\in I_{t}^{\star}:i\neq j}\lambda_{i,j}\left(A_{i,j}(h)+\frac{1}{2L}C_{i,j}\right)=Z\\ &Z\succeq 0\\ &v,w\geq 0,\;\lambda_{i,j}\geq 0,\quad\forall i,j\in I_{t}^{\star}:i\neq j\ .\end{cases}$ (3.3)
Although it is not needed for our analysis, equality holds here as well (i.e., strong duality holds) due to [2, Theorem 6].
@@ 3.2 An SDP Feasibility Certificate that implies Straightforwardness
where $m_{h}\colon\mathbb{R}^{(t+2)\times(t+2)}\rightarrow\mathbb{R}^{t+1}$ and $M_{h}\colon\mathbb{R}^{(t+2)\times(t+2)}\rightarrow\mathbb{R}^{(t+1)\times(t+1)}$ are linear functions. Certifying $p_{1,1}(\delta)\leq\delta-\sum_{i=0}^{t-1}(h_{i}-\epsilon)\delta^{2}$ for fixed $\delta$ then follows by showing the following spectral set is nonempty
@@
$\begin{bmatrix}\sum_{i=0}^{t-1}(h_{i}+\epsilon)\delta^{2}&\delta m_{h}(\gamma)^{T}\\ \delta m_{h}(\gamma)&M_{h}(\lambda)+\delta M_{h}(\gamma)\end{bmatrix}\succeq 0\ .$
@@
$\begin{bmatrix}\sum_{i=0}^{t-1}(h_{i}+\epsilon)&m_{h}(\gamma)^{T}\\ m_{h}(\gamma)&M_{h}(\lambda)+\delta M_{h}(\gamma)\end{bmatrix}\succeq 0\ .$
@@
Note computing a member of $\mathcal{S}_{h,\epsilon,\Delta}$ does not correspond to solving a particular performance estimation problem. Instead, each member provides a “line segment” of solutions to a series of the performance estimation problems $p_{L,D}(\delta)$ for an interval of possible $\delta$ values.
3.3 Certificates of Straightforwardness Proving Guarantees in Table 1